OpenStudy (anonymous):

The table shows the proof of the relationship between the slopes of two perpendicular lines. What is the missing statement in step 2?

OpenStudy (anonymous):

1. is similar to . given 2. property of similar triangles 3. = property of proportion 4. slope of = slope of = definition of slope 5. slope of × slope of = multiplying the slopes 6. (This statement is intentionally left blank.) Substitution Property of Equality 7. slope of × slope of = -1 simplifying the right side

OpenStudy (anonymous):

couldnt get all of it but tried my best

jimthompson5910 (jim_thompson5910):

a screenshot is best because we can see the full problem (even the missing symbols)

OpenStudy (anonymous):

1. AC←→⊥CD←→ (view diagram) ΔABC is similar to ΔCED. given 2. (This statement is intentionally left blank.) property of similar triangles 3. ABBC = ECDE property of proportion 4. slope of AC←→ = -ABBC slope of DC←→ = DEAC definition of slope 5. slope of AC←→ × slope of DC←→ = -ABBC×DEEC multiplying the slopes 6. Substitution Property of Equality 7. slope of AC←→ × slope of DC←→ = -1 simplifying the right side

OpenStudy (anonymous):

sorry here it is

jimthompson5910 (jim_thompson5910):

well step 3 has AB*BC = EC*DE, or so it seems the step before that would be some proportion and cross multiplying would lead to step 3

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

not really but u can explain it in detail. mabye ill get it

jimthompson5910 (jim_thompson5910):

let's say you had a/b = c/d

jimthompson5910 (jim_thompson5910):

if you cross multiplied, you'd get a*d = b*c agreed?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

so what proportion could be set up to lead to AB*BC = EC*DE ?

OpenStudy (anonymous):

they seem to look like this but in fraction AB/BC;. something like that.

OpenStudy (anonymous):

AB/DE, and EC/DE

jimthompson5910 (jim_thompson5910):

more like AB/EC = DE/BC. There are other possibilities